The official record for swimming the English Channel, a distance of 18 miles, is held by Penny Dean, who completed her swim in 7 hrs 40 minutes What was her average speed, in miles per hour?

Answer by checkley71(8403)

distance=rate*time
18=r*420+40
18=r*460
r=18/460
r=.039 miles/minute or 60*.039=2.35 miles/hour.